OpenStudy (anonymous):

Identify the discontinuities, graph, and find the zeros of f(x) = 4x^2 - 36x / x - 9?

OpenStudy (anonymous):

Can someone help??

OpenStudy (anonymous):

evidently it will not be continuous at \(x=9\) because you cannot divide by zero

OpenStudy (anonymous):

To find the 0's of the function set f(x) = 0 and solve for x

OpenStudy (anonymous):

\[\frac{4x^2-36x}{x-9}=\frac{4x(x-9)}{x-9}=4x\] if \(x\neq 9\)

OpenStudy (anonymous):

Ok so discounity is 9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and it will be 0 if \(x=0\)

OpenStudy (anonymous):

I got the 4x but what would that be for?

OpenStudy (anonymous):

the function is really just \(f(x)=4x\) which is a line with slope 4 through the origin but it is undefined if \(x=9\)

OpenStudy (anonymous):

Ok so discontinuities is 9 and the zeros you find by putting f(0) and solving it like that?

OpenStudy (anonymous):

evidently \(4x=0\) if \(x=0\) so the answer to "find the zeros" is \(x=0\)

OpenStudy (anonymous):

you do not find the zeros by finding \(f(0)\) but rather by setting \(f(x)=0\) and solving for \(x\) it just so happens that in this case \(x=0\) is also the answer

OpenStudy (anonymous):

I still a little stuck

OpenStudy (anonymous):

ok lets answer the questions one by one

OpenStudy (anonymous):

Ok thanks :)!

OpenStudy (anonymous):

Identify the discontinuities answer: it is discontinuous at \(x=9\) because you cannot divide by 0

OpenStudy (anonymous):

That part I got :)

OpenStudy (anonymous):

The rest just confuses me :(

OpenStudy (anonymous):

graph answer: the graph is identical to the graph of the line \(y=4x\) a line with slope \(4\) through the origin the only difference is that at \((9,36)\) since the original function is not defined at \(x=9\) you have a hole

OpenStudy (anonymous):

So you times 4 by the 9 and got 36 for y, right?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes, exactly

OpenStudy (anonymous):

So I so far get 9 is the discounity and by zeros are (9,36) is that correct?

OpenStudy (anonymous):

find the zeros answer : if \(x=0\) then \(f(x)=0\)

OpenStudy (anonymous):

no, the discontinuity is at \((9,36)\) the zero is at \(x=0\) because that is where the line crosses the \(x\) axis

OpenStudy (anonymous):

they are two separate concepts

OpenStudy (anonymous):

So wherever the line crosses the x axis is the zero?

OpenStudy (anonymous):

yes it is also the solution or the solutions to \(f(x)=0\)

OpenStudy (anonymous):

So I have to graph to find the zero?

OpenStudy (anonymous):

no not really you can set \(4x=0\) and solve for \(x\) (in your head)

OpenStudy (anonymous):

Ohh I see what you did

OpenStudy (anonymous):

So the zero=0

OpenStudy (anonymous):

and discounity is (9, 36)

OpenStudy (anonymous):

yes or you could just say the discontinuity is at \(x=9\) either way

OpenStudy (anonymous):

Am I missing anything else?

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

Thanks! for your help! :) :)

OpenStudy (anonymous):

yw